# Two-Player Zero-Sum Games with Bandit Feedback

### v2

@@ Appendix D Proof of Theorem 4
$\mathbb{P}(E_{1})$ $=\mathbb{P}(\hat{A}(i^{\prime},j)-\hat{A}(i,j)\leq\varepsilon_{t},\forall i^{\prime}\in S_{x})$ (66)
$\leq\mathbb{P}(\hat{A}(i^{\prime},j)-\hat{A}(i,j)\leq\varepsilon_{t})$ (67)
$\leq\exp{\Bigg(-\frac{\frac{4\sigma^{2}}{k_{t}}\ln{\Big(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\Big)}}{2(\frac{\sigma^{2}}{k_{i^{\prime}j,t}}+\frac{\sigma^{2}}{k_{ij,t}})}\Bigg)}$ (68)
$\leq\exp{\Bigg(-\frac{\frac{4\sigma^{2}}{k_{t}}\ln{\Big(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\Big)}}{2\Big(\frac{\sigma^{2}}{(k_{t}+k_{t}\hat{P}_{i^{\prime}j,t})}+\frac{\sigma^{2}}{(k_{t}+k_{t}\hat{P}_{ij,t})}\Big)}\Bigg)}$ (69)
$\leq\exp{\Bigg(-\frac{\frac{4\sigma^{2}}{k_{t}}\ln{\Big(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\Big)}}{2\sigma^{2}\Big(\frac{(\hat{P}_{i^{\prime}j,t}+\hat{P}_{ij,t}+2)}{k_{t}(1+\hat{P}_{i^{\prime}j,t})(1+\hat{P}_{ij,t})}\Big)}\Bigg)}$ (70)
$=\exp{\Bigg(-\frac{2(1+\hat{P}_{i^{\prime}j,t})(1+\hat{P}_{ij,t})}{(\hat{P}_{i^{\prime}j,t}+\hat{P}_{ij,t}+2)}\ln{\Big(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\Big)}\Bigg)}$ (71)
$\leq\exp{\Bigg(-\ln{\Big(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\Big)}}\Bigg)$ (72)
$=\frac{16\sigma^{2}}{\hat{\Delta}_{t}^{2}T}$ (73)
Since the term $\frac{2(1+\hat{P}_{i^{\prime}j,t})(1+\hat{P}_{ij,t})}{(\hat{P}_{i^{\prime}j,t}+\hat{P}_{ij,t}+2)}$ is greater than one, the inequality (72) follows. In addition, since we obtain the same result for $\mathbb{P}(E_{2})$ as $\mathbb{P}(E_{2})$, which leads to the same probability of keeping an action pair as in the proof of Theorem 2 and also the same results for Case 1, Case 3.
@@
$k_{t_{ij}}=\Big\lceil\frac{16\sigma^{2}}{\hat{\Delta}_{t_{ij}}^{2}}\ln{(\frac{\hat{\Delta}_{t_{ij}}^{2}T}{16\sigma^{2}})}\Big\rceil+\Big\lceil\frac{16\sigma^{2}\hat{P}_{ij,t_{ij}}}{\hat{\Delta}_{t_{ij}}^{2}}\ln{(\frac{\hat{\Delta}_{t_{ij}}^{2}T}{16\sigma^{2}})}\Big\rceil\leq 2\ \Big\lceil\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}\Big\rceil$

### v3

@@ Appendix D Proof of Theorem 4
$\mathbb{P}(E_{1}\cap E_{2})$ $\leq\min\{\mathbb{P}(E_{1}),\mathbb{P}(E_{2})\}$ (79)
$\leq\min\Bigg\{\exp{\Bigg(-\frac{(\Delta_{ij}^{\text{max}}-\varepsilon_{t})^{2}}{2(\frac{\sigma^{2}}{k_{i^{\prime}j,t}}+\frac{\sigma^{2}}{k_{ij,t}})}\Bigg)},\exp{\Bigg(-\frac{(\Delta_{ij}^{\text{min}}-\varepsilon_{t})^{2}}{2(\frac{\sigma^{2}}{k_{ij^{\prime},t}}+\frac{\sigma^{2}}{k_{ij,t}})}\Bigg)}\Bigg\}$ (80)
$\leq\min\Bigg\{\exp{\Bigg(-\frac{(\Delta_{ij}^{\text{max}}-\varepsilon_{t})^{2}}{2\Big(\frac{\sigma^{2}}{(k_{t}+k_{t}\hat{P}_{i^{\prime}j,t})}+\frac{\sigma^{2}}{(k_{t}+k_{t}\hat{P}_{ij,t})}\Big)}\Bigg)},\exp{\Bigg(-\frac{(\Delta_{ij}^{\text{min}}-\varepsilon_{t})^{2}}{2\Big(\frac{\sigma^{2}}{(k_{t}+k_{t}\hat{P}_{ij^{\prime},t})}+\frac{\sigma^{2}}{(k_{t}+k_{t}\hat{P}_{ij,t})}\Big)}\Bigg)}\Bigg\}$ (81)
$\leq\min\Bigg\{\exp{\Bigg(-\frac{(\Delta_{ij}^{\text{max}}-\varepsilon_{t})^{2}}{2\sigma^{2}\Big(\frac{(\hat{P}_{i^{\prime}j,t}+\hat{P}_{ij,t}+2)}{k_{t}(1+\hat{P}_{i^{\prime}j,t})(1+\hat{P}_{ij,t})}\Big)}\Bigg)},\exp{\Bigg(-\frac{(\Delta_{ij}^{\text{min}}-\varepsilon_{t})^{2}}{2\sigma^{2}\Big(\frac{(\hat{P}_{ij^{\prime},t}+\hat{P}_{ij,t}+2)}{k_{t}(1+\hat{P}_{ij^{\prime},t})(1+\hat{P}_{ij,t})}\Big)}\Bigg)}\Bigg\}$ (82)
$\leq\min\Bigg\{\exp{\Bigg(-\frac{k_{t}(\Delta_{ij}^{\text{max}}-\varepsilon_{t})^{2}}{4\sigma^{2}}\Bigg)},\exp{\Bigg(-\frac{k_{t}(\Delta_{ij}^{\text{min}}-\varepsilon_{t})^{2}}{4\sigma^{2}}\Bigg)}\Bigg\}$ (83)
$\leq\frac{16\sigma^{2}}{\hat{\Delta}_{t}^{2}T}.$ (84)
Since the terms $\frac{(\hat{P}_{i^{\prime}j,t}+\hat{P}_{ij,t}+2)}{(1+\hat{P}_{i^{\prime}j,t})(1+\hat{P}_{ij,t})}$ and $\frac{(\hat{P}_{ij^{\prime},t}+\hat{P}_{ij,t}+2)}{(1+\hat{P}_{ij^{\prime},t})(1+\hat{P}_{ij,t})}$ are obviously less than 2 since $0\leq\hat{P}_{ij,t}\leq 1$ for any $(i,j)\in S_{A}$, the inequality (83) follows. Then, it leads to the same probability of keeping an action pair as in the proof of Theorem 2 and also to the same results for Case 1, Case 3.
@@
$k_{t_{ij}}=\Big\lceil\frac{16\sigma^{2}}{\hat{\Delta}_{t_{ij}}^{2}}\ln{(\frac{\hat{\Delta}_{t_{ij}}^{2}T}{16\sigma^{2}})}\Big\rceil+\Big\lceil\frac{16\sigma^{2}\hat{P}_{ij,t_{ij}}}{\hat{\Delta}_{t_{ij}}^{2}}\ln{(\frac{\hat{\Delta}_{t_{ij}}^{2}T}{16\sigma^{2}})}\Big\rceil\leq 2\ \Big\lceil\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}\Big\rceil$ (85)